# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ 3 LAGRANGIAN RELAXATION OF THE LOWER BOUND
In this section, we discuss the Lagrangian relaxation of the lower bound and its properties, necessary to design a correct and feasible $r$-good pure exploration algorithm under unknown linear constraints. We require two structural assumptions.
Assumption 1 (Assumptions on means, policy, and constraints).
(a) The mean vector ${\boldsymbol{\mu}}$ belongs to a bounded subset $\mathcal{D}$ of $\mathbb{R}^{{K}}$. (b) There exists a unique optimal feasibly policy (Equation (1)). (c) For the true constraint $\mathbf{A}$, there exists a non-zero slack vector $\Gamma$, such that $\max_{\boldsymbol{\pi}\in\Delta_{K}}(-\mathbf{A}\boldsymbol{\pi})=\Gamma.$
We impose the unique optimal and feasible policy assumption following Carlsson et al., (2024) to ensure that the solution of Equation (1) is a unique extreme point of the polytope $\mathcal{F}$. The assumption on slack is analogous to assuming the existence of a safe-arm (Pacchiano et al.,, 2020), or that of Slater’s condition for the constraint optimisation problem (Liu et al.,, 2021). Standing on these assumptions, we prove that $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$ is unique, i.e $\boldsymbol{\pi}^{\star}_{\hat{\mathcal{F}}}$ is an extreme point in the polytope $\hat{\mathcal{F}}$.
3.1 Information Acquisition: Estimating Constraints
The agent acquires information at every step $t\in\mathbb{N}$ by sampling an action $\mathbf{a}_{t}\sim\boldsymbol{\omega}_{t}$. $\boldsymbol{\omega}_{t}\in\Delta_{K}$ is called the allocation policy. This yields a noisy reward $r_{t}\in\mathbb{R}$ and cost vector $\mathbf{A}\mathbf{a}_{t}\in\mathbb{R}^{d}$. As the arms are independent, we represent the $a$-th arm as the $a$-th basis of $\mathbb{R}^{K}$. Thus, using the observations obtained till $t$, we estimate the mean vector as $\hat{{\boldsymbol{\mu}}}_{t}\mathrel{\triangleq}\Sigma_{t}^{-1}\left(\sum_{s=1}^{t-1}r_{s}\mathbf{a}_{s}\right)$. Here, $\Sigma_{t}\mathrel{\triangleq}v\mathbf{I}+\sum_{s=1}^{t}\mathbf{a}_{s}\mathbf{a}_{s}^{\top}$ for any $v>0$, is the Gram matrix or the design matrix. Similarly, the estimate of the $i$-th row of the constraint matrix is $\hat{\mathbf{A}}_{t}^{i}\mathrel{\triangleq}\Sigma_{t}^{-1}\left(\sum_{s=1}^{t-1}\mathbf{A}^{i,\mathbf{a}_{s}}\mathbf{a}_{s}\right)$. But naively using $\hat{\mathbf{A}}_{t}$ to define the feasible policy set does not ensure that for any $t$, the estimated feasible set $\hat{\mathcal{F}}$ is a superset of $\mathcal{F}$. Hence, we define a confidence ellipsoid around $\hat{\mathbf{A}}_{t}$ that includes $\mathbf{A}$ with probability at least $1-\delta$, and construct a pessimistic estimate of $\mathbf{A}$. Formally, the confidence ellipsoid is
$\mathcal{C}_{t}\mathrel{\triangleq}\big{\{}\mathbf{A}^{\prime}\in\mathbb{R}^{d\times{K}}\mid\|\mathbf{A}^{\prime i}-\hat{\mathbf{A}}_{t}^{i}\|_{\Sigma_{t}}\leq f(t,\delta)\forall i\in[d]\big{\}},$ (2)
@@ 3.1 Information Acquisition: Estimating Constraints
Lemma 1 (Pessimistic feasible sets).
If we construct the pessimistic feasible policy set at any time $t\in\mathbb{N}$ as
^$\hat{\mathcal{F}}_{t}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}:\min_{\mathbf{A}^{\prime}\in\mathcal{C}_{t}}\mathbf{A}^{\prime}\boldsymbol{\pi}\leq\boldsymbol{0}\}\,,\vspace*{-1.8em}$ (3)
we observe that ^$\mathcal{F}\subseteq\hat{\mathcal{F}}_{t}$ with probability $1-\delta$.
We denote the $\mathbf{A}^{\prime}$ achieving the above minimum as $\tilde{\mathbf{A}}_{t}$. In Figure 1, we visualise this result using the numerical values obtained from our algorithms. We observe that as we acquire more samples, our estimated feasible policy set ^$\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$.
Remark. To ensure that the true optimal and feasible policy $\boldsymbol{\pi}^{\star}_{\mathcal{F}}\in\hat{\mathcal{F}}_{t}$ is inside the estimated feasible policy set, and ^$\hat{\mathcal{F}}_{t}\neq\emptyset$ for any $t$, we use the pessimistic estimate of the feasible set ^$\hat{\mathcal{F}}_{t}$ ensuring ^$\mathcal{F}\subseteq\hat{\mathcal{F}}_{t}$ with high probability. Our construction of the pessimistic estimates of feasible set resonate with the spirit of optimistic-pessimistic algorithms for regret-minimisation under constraints (Pacchiano et al.,, 2020; Liu et al.,, 2021; Chen et al., 2022b, ; Pacchiano et al.,, 2024).
Figure 1: Convergence of the pessimistic feasible set and optimal policy.
3.2 Lagrangian Relaxation with Estimated Constraints
Search for the optimal policy is essentially a linear programming problem when we know the mean vector ${\boldsymbol{\mu}}$ and a constraint matrix $\mathbf{A}$. The challenge in bandit is to identify them from sequential feedback, i.e. to differentiate ${\boldsymbol{\mu}}$ from the other confusing instances in the same family of distributions. These are called the alternative instances. The strategy is to gather enough statistical evidence to rule out all such confusing instances, specifically the one that has minimum KL-divergence from ${\boldsymbol{\mu}}$ as observed under the allocation policy $\boldsymbol{\omega}$ (Garivier and Kaufmann,, 2016). This intuition has led to the lower bound of Carlsson et al., (2024) stating that the expected stopping time of any $(1-\delta)$-correct and always-feasible algorithm satisfies
@@ 3.2 Lagrangian Relaxation with Estimated Constraints
if $\mathbf{A}$ is known. $T_{\mathcal{F},r}({\boldsymbol{\mu}})$ is called the characteristics time. Its reciprocal is a max-min optimisation problem over the set of alternative instances $\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})\mathrel{\triangleq}\big{\{}{\boldsymbol{\lambda}}\in\mathcal{D}\mid\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}-r>\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}\mathrel{\triangleq}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}^{\star}\big{\}}$, i.e.
@@
$T^{-1}_{\mathcal{F},r}({\boldsymbol{\mu}})$ $\mathrel{\triangleq}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\pi\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a}\right)$
$\mathrel{\triangleq}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\pi\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)\,.$ (5)
@@
$\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})$, referred as the Alt-set, is the set of all bandit instances whose mean vectors are in a bounded subset $\mathcal{D}\in\mathbb{R}^{K}$ but the optimal policy is different than that of ${\boldsymbol{\mu}}\in\mathcal{D}$. Now, we inspect the change in this lower bound at any step $t>0$, when we only have access to a pessimistic estimate ^$\hat{\mathcal{F}}_{t}$ and the confidence ellipsoid $\mathcal{C}_{t}$ but do not know $\mathcal{F}$. For brevity, we exclude $t$ from the subscripts for where it is clear from the context. Now, we observe that the Alt-set given $\hat{\mathcal{F}}$ is
$\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})\mathrel{\triangleq}\big{\{}{\boldsymbol{\lambda}}\in\mathcal{D}\mid\max_{\boldsymbol{\pi}\in\hat{\mathcal{F}}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}-r>{\boldsymbol{\lambda}}^{\top}\hat{\boldsymbol{\pi}}^{\star}\big{\}}$ (6)
@@
Since $\mathcal{F}\subseteq\hat{\mathcal{F}}$, we observe that $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})\subseteq\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})$. Now, we are ready define to Lagrangian relaxation of the lower bound, i.e.
@@
$T^{-1}_{\mathcal{F},r}({\boldsymbol{\mu}})$ $\mathrel{\triangleq}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)$
$\leq\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{\mathbf{A}^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\mathbf{A}^{\prime}\boldsymbol{\omega}$ (7)
@@
We denote this Lagrangian relaxation of the characteristic time with $\hat{\mathcal{F}}$ as $T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})$. For non-negative Lagrange multipliers $\boldsymbol{l}\in\mathbb{R}_{+}^{d}$, the first inequality is true due to the existence of a slack for the true constraints $\mathbf{A}$. The second inequality is due to the pessimistic choice of the estimated constraint. Equation (7) shows that the reciprocal of the Lagrangian relaxation, $T_{\hat{\mathcal{F}},r}({\boldsymbol{\mu}})$, serves as a upper bound on the characteristic time $T_{\mathcal{F}}({\boldsymbol{\mu}})$ for known constraints (Carlsson et al.,, 2024).
@@
The optimisation problem in Equation (7) satisfies
$\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{\mathbf{A}^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}{\mathbf{A}^{\prime}}\boldsymbol{\omega}$
$=$ $\sup_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\,.$ (8)
@@
Here, $\mathcal{L}\mathrel{\triangleq}\{\boldsymbol{l}\in\mathbb{R}_{+}^{d}\mid 0\leq\|{\boldsymbol{l}}\|_{1}\leq\frac{1}{\gamma}\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\}$, where $\gamma\mathrel{\triangleq}\min_{i\in[1,{d}]}\{-\tilde{\mathbf{A}}^{i}\boldsymbol{\omega}^{*}\}$, i.e. the minimum slack w.r.t. the optimal allocation. (Proof is in Appendix C.)
@@
Hereafter, we use the RHS of Eq. (1) as $T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})$. Theorem 1 provides a hypercube to search for the Lagrangian multipliers, which is a linear programming problem.
@@
Remark: Connections with Lagrangian-based Methods in Bandits. Regret minimisation literature leverages Lagrangian-based optimistic-pessimistic methods (Tirinzoni et al.,, 2020; Slivkins et al.,, 2023) to obtain both the sub-linear regret and constraint violation guarantees (Liu et al.,, 2021; Bernasconi et al.,, 2024). Our proposed algorithm LAGEX (Algorithm 2) is a prime example where the "self-boundedness" of the dual variables results in tighter constraint violation guarantees (Figure 7 and 7).
Inner Optimisation Problem. Now, we peel the layers of the optimisation problem in Eq. (1) and focus on obtaining
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\triangleq\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{T}\tilde{\mathbf{A}}\boldsymbol{\omega}\,.$
@@
For known constraints and $r=0$, Carlsson et al., (2024) has leveraged results from convex analysis (Boyd and Vandenberghe,, 2004) to show that the most confusing instance for ${\boldsymbol{\mu}}$ lie in the boundary of the normal cone $\Lambda_{\mathcal{F}}({\boldsymbol{\mu}},\boldsymbol{\pi})^{C}$ spanned by the active constraints $\mathbf{A}_{\boldsymbol{\pi}^{\star}_{\mathcal{F}}}$ for $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$. $\mathbf{A}_{\boldsymbol{\pi}^{\star}_{\mathcal{F}}}$ is a sub-matrix of $\mathbf{A}$ consisting at least $K$ linearly independent rows. This is called the projection lemma. Specifically, $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F}\mid r=0)=\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{0}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\mathcal{F}}(\boldsymbol{\pi})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{T}(\boldsymbol{\pi}-\boldsymbol{\pi}^{{}^{\prime}})=0}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)$. In our setting, we are sequentially estimating both the mean vectors and the constraints, and thus, the normal cone. Now, we derive the projection lemma for our pessimistic feasible set and $r\neq 0$.

### v2

@@ 3 Lagrangian Relaxation of the Lower Bound
Now, we derive Lagrangian relaxation of lower bound and its properties under a structural assumption.
Assumption 1 (Structures of means, policy, and constraints).
(a) The mean vector $\boldsymbol{\mu}$} is in a bounded subset $\mathcal{D}$ of $\mathbb{R}^{{K}}$. (b) There exists a unique optimal feasibly policy (Eq. (1)). (c) The true constraint $A$ yields a non-zero slack vector $\Gamma$: $\max_{\boldsymbol{\pi}\in\Delta_{K}}(-A\boldsymbol{\pi})=\Gamma.$
We impose the unique optimal and feasible policy assumption following Carlsson et al., (2024). The assumption on slack is analogous to existence of a safe-arm (Pacchiano et al.,, 2020), or Slater’s condition for the constraint optimisation problem (Liu et al.,, 2021).
3.1 Information Acquisition: Estimate Constraints
The agent acquires information at every step $t\in\mathbb{N}$ by sampling an action $\mathbf{a}_{t}\sim\boldsymbol{\omega}_{t}$. As the arms are independent, we represent the $a$-th arm as the $a$-th basis in $\mathbb{R}^{K}$, denoted by $\mathbf{a}\in\mathbb{R}^{K}$. $\boldsymbol{\omega}_{t}\in\Delta_{K}$ is called the allocation policy. As shown in Algorithm 1, pulling the arm $a_{t}$ yields a noisy reward $r_{t}\in\mathbb{R}$ and cost vector $\mathbf{c}_{t}\in\mathbb{R}^{d}$. The cost vector $\mathbf{c}_{t}\mathrel{\triangleq}A\mathbf{a}_{t}+\boldsymbol{\eta}_{t}$, where $\boldsymbol{\eta}_{t}$ is an independent and identically distributed noise vector whose each component uis generated from a $1$-sub-Gaussian distribution with mean zero.
Thus, using the observations obtained till $t$, we estimate the mean vector as $\hat{{\boldsymbol{\mu}}}_{t}\mathrel{\triangleq}\Sigma_{t}^{-1}\left(\sum_{s=1}^{t-1}r_{s}\mathbf{a}_{s}\right)$. Here, $\Sigma_{t}\mathrel{\triangleq}I+\sum_{s=1}^{\top}\mathbf{a}_{s}\mathbf{a}_{s}^{\top}$, is the Gram matrix or the design matrix. Similarly, the estimate of the $i$-th row of the constraint matrix is $\hat{A}_{t}^{i}\mathrel{\triangleq}\Sigma_{t}^{-1}\left(\sum_{s=1}^{t-1}A^{i,\mathbf{a}_{s}}\mathbf{a}_{s}\right)$. But naïvely using $\hat{A}_{t}$ to define the feasible policy set does not ensure that for any $t$, the estimated feasible set $\hat{\mathcal{F}}$ is a superset of $\mathcal{F}$. Hence, we define a confidence ellipsoid around $\hat{A}_{t}$ that includes $A$ with probability at least $1-\delta$, and construct a optimistic estimate of $A$. Formally, the confidence ellipsoid is
$\hskip-10.00002pt\mathcal{C}_{t}\mathrel{\triangleq}\big\{A^{\prime}\in\mathbb{R}^{d\times{K}}\mid\|A^{\prime i}-\hat{A}_{t}^{i}\|_{\Sigma_{t}}\leq f(t,\delta)\forall i\in[d]\big\},$ (2)
@@ 3.1 Information Acquisition: Estimate Constraints
Lemma 1 (Optimistic feasible sets).
At any time $t\in\mathbb{N}$, we construct the optimistic feasible policy set such that with probability $1-\delta$, $\hat{\mathcal{F}}_{t}\mathrel{\triangleq}\{\boldsymbol{\pi}\in\Delta_{K}:\min_{A^{\prime}\in\mathcal{C}_{t}}A^{\prime}\boldsymbol{\pi}\leq\boldsymbol{0}\}$, satisfies $\mathcal{F}\subseteq\hat{\mathcal{F}}_{t}$, where $\tilde{A}_{t}\triangleq\operatorname*{arg\,min}_{A^{\prime}\in\mathcal{C}_{t}}A^{\prime}\boldsymbol{\pi}$.
Figure 3 plots this result using the numerical values obtained from our algorithms. Note that as we acquire more samples, estimated feasible policy set $\hat{\mathcal{F}}_{t}\rightarrow\mathcal{F}$.
Remark 1.
To ensure that the true optimal and feasible policy $\boldsymbol{\pi}^{\star}_{\mathcal{F}}\in\hat{\mathcal{F}}_{t}$ is inside the estimated feasible policy set, and $\hat{\mathcal{F}}_{t}\neq\emptyset$ for any $t$, we use the optimistic estimate of the feasible set $\hat{\mathcal{F}}_{t}$ ensuring $\mathcal{F}\subseteq\hat{\mathcal{F}}_{t}$ with high probability. Our construction of the optimistic estimates of feasible set resonate with the optimistic-pessimistic algorithms for regret-minimisation under constraints (Pacchiano et al.,, 2020, 2024; Liu et al.,, 2021; Chen et al., 2022b, ). Note that, similar confidence bounds for estimating constraints are used to test feasibility of linear programs (Gangrade et al.,, 2024).
3.2 Relaxation with Estimated Constraints
Search for the optimal policy is essentially a linear programming problem when we know the mean vector $\boldsymbol{\mu}$} and a constraint matrix $A$. The challenge in bandit is to identify them from sequential feedback, i.e. to differentiate $\boldsymbol{\mu}$} from the other confusing instances in the same family of distributions. These are called the alternative instances. The strategy is to gather enough statistical evidence to rule out all such confusing instances, specifically the one that has minimum KL-divergence from $\boldsymbol{\mu}$} as observed under the allocation policy $\boldsymbol{\omega}$ (Garivier and Kaufmann,, 2016). This intuition has led to the lower bound of Carlsson et al., (2024)– the expected stopping time of any $(1-\delta)$-correct and always-feasible algorithm satisfies
@@ 3.2 Relaxation with Estimated Constraints
if $A$ is known. $T_{\mathcal{F},r}({\boldsymbol{\mu}})$ is called the characteristics time. Its reciprocal is a max-min optimisation problem over the set of alternative instances $\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})\mathrel{\triangleq}\big\{{\boldsymbol{\lambda}}\in\mathcal{D}\mid\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}-r>{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}^{\star}\big\}$, where $\boldsymbol{\pi}^{\star}\in\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$, i.e.
@@
$T^{-1}_{\mathcal{F},r}({\boldsymbol{\mu}})\mathrel{\triangleq}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\pi\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d\left({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a}\right)\mathrel{\triangleq}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\pi\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)\,.$ (4)
@@
$\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})$, referred as the Alt-set, is the set of all bandit instances whose mean vectors are in a bounded subset $\mathcal{D}\in\mathbb{R}^{K}$ but the optimal policy is different than that of ${\boldsymbol{\mu}}\in\mathcal{D}$. Now, we inspect the change in this lower bound at any step $t>0$, when we only have access to a optimisti estimate $\hat{\mathcal{F}}_{t}$ and the confidence ellipsoid $\mathcal{C}_{t}$ but do not know $\mathcal{F}$. For brevity, we exclude $t$ from the subscripts for where it is clear from the context. The Alt-set given $\hat{\mathcal{F}}$ is defined as
$\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})\mathrel{\triangleq}\big\{{\boldsymbol{\lambda}}\in\mathcal{D}\mid\max_{\boldsymbol{\pi}\in\hat{\mathcal{F}}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}-r>{\boldsymbol{\lambda}}^{\top}\hat{\boldsymbol{\pi}}^{\star}\big\},$ (5)
@@
where $\hat{\boldsymbol{\pi}}^{\star}\in\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}{\boldsymbol{\mu}}^{\top}\boldsymbol{\pi}$. Since $\mathcal{F}\subseteq\hat{\mathcal{F}}$, we observe that $\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})\subseteq\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})$ (Figure 3). Now, we define Lagrangian relaxation of the lower bound, i.e.
@@
$\hskip-60.00009ptT^{-1}_{\mathcal{F},r}({\boldsymbol{\mu}})\mathrel{\triangleq}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)\leq\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{A^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}A^{\prime}\boldsymbol{\omega}$ (6)
@@
We denote this Lagrangian relaxation of the characteristic time with $\hat{\mathcal{F}}$ as $T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})$. For non-negative Lagrange multipliers $\boldsymbol{l}\in\mathbb{R}_{+}^{d}$, the first inequality is true due to the existence of a slack for the true constraints $A$. The inequality holds due to the optimistic choice of the estimated constraint. Equation (6) shows that the reciprocal of the Lagrangian relaxation, $T_{\hat{\mathcal{F}},r}({\boldsymbol{\mu}})$, serves as a upper bound on the characteristic time $T_{\mathcal{F}}({\boldsymbol{\mu}})$ for known constraints (Carlsson et al.,, 2024).
@@
We formalise the strong duality result and self-bounding property of the Lagrangian multiplier of the relaxation in Equation (6) in Theorem 1 below.
@@
The optimisation problem in Equation (6) satisfies $\inf_{\boldsymbol{l}\in\mathbb{R}_{+}^{d}}\min_{A^{\prime}\in\mathcal{C}}\sup_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}{A^{\prime}}\boldsymbol{\omega}=\sup_{\boldsymbol{\omega}\in\Delta_{K}}\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\,.$ (7) Here, $\mathcal{L}\mathrel{\triangleq}\{\boldsymbol{l}\in\mathbb{R}_{+}^{d}\mid 0\leq\|{\boldsymbol{l}}\|_{1}\leq\frac{1}{\gamma}\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F})\}$, where $\gamma\mathrel{\triangleq}\min_{i\in[1,{d}]}\{-\tilde{A}^{i}\hat{\boldsymbol{\pi}^{\star}}\}$, i.e. the minimum slack w.r.t. the estimated optimal feasible policy.
@@
Detailed proof is in Appendix C. Hereafter, we use the RHS of Eq. (7) as $T_{\hat{\mathcal{F}},r}^{-1}({\boldsymbol{\mu}})$. Theorem 1 provides a hypercube to search for the Lagrangian multipliers, which is a linear programming problem.
Figure 2: Convergence of the optimistic feasible set and optimal policy.
Figure 3: Normal cones over time.
Remark 2 (Connections with Lagrangian-based Methods in Bandits.).
@@
Regret minimisation literature leverages Lagrangian-based optimistic-pessimistic methods (Tirinzoni et al.,, 2020; Slivkins et al.,, 2023) to obtain both the sub-linear regret and constraint violation guarantees (Liu et al.,, 2021; Bernasconi et al.,, 2024). Our proposed algorithm LAGEX (Algorithm 3) is a prime example where the “self-boundedness” of the dual variables results in tighter constraint violation guarantees (Figure 9 and 9).
I. The Inner Optimisation Problem. Now, we peel the layers of the optimisation problem in Eq. (7) and focus on obtaining
@@
$\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\hat{\mathcal{F}})\triangleq\min_{\boldsymbol{l}\in\mathcal{L}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right)-\boldsymbol{l}^{\top}\tilde{A}\boldsymbol{\omega}\,.$
@@
For known constraints and $r=0$, Carlsson et al., (2024) has leveraged results from convex analysis (Boyd and Vandenberghe,, 2004) to show that the most confusing instance for $\boldsymbol{\mu}$} lie in the boundary of the normal cone $\Lambda_{\mathcal{F}}({\boldsymbol{\mu}})^{C}$ spanned by the active constraints $A_{\boldsymbol{\pi}^{\star}_{\mathcal{F}}}$ for $\boldsymbol{\pi}^{\star}_{\mathcal{F}}$. $A_{\boldsymbol{\pi}^{\star}_{\mathcal{F}}}$ is a sub-matrix of $A$ consisting at least $K$ linearly independent rows. This is called the projection lemma. Specifically, $\mathcal{D}(\boldsymbol{\omega},{\boldsymbol{\mu}},\mathcal{F}\mid r=0)=\max_{\boldsymbol{\pi}\in\Pi_{\mathcal{F}}^{0}}\min_{\boldsymbol{\pi}^{{}^{\prime}}\in\nu_{\mathcal{F}}(\boldsymbol{\pi})}\min_{{\boldsymbol{\lambda}}:{\boldsymbol{\lambda}}^{\top}(\boldsymbol{\pi}-\boldsymbol{\pi}^{{}^{\prime}})=0}\boldsymbol{\omega}^{\top}d\left({\boldsymbol{\mu}},{\boldsymbol{\lambda}}\right).$
In our setting, we are sequentially estimating both the mean vectors and the constraints, and thus, the normal cone. Now, we derive the projection lemma for our optimisti feasible set and $r\neq 0$.